OpenStudy (anonymous):

Simplify the given expression: (1 + 3i) + (2 − 5i) −1 − 2i 3 − 2i 4 − 3i 5 + 6i

OpenStudy (anonymous):

given expression is 1 + 3i +2 -5i = answer, you understand why?

OpenStudy (anonymous):

No , whats the answer lol?

OpenStudy (anonymous):

@BoneFreeze

OpenStudy (anonymous):

??

OpenStudy (anonymous):

sole purpose of this site is that you learn through engagement with problem solving and not that you get answers through using other people's manifestation with getting medals :) look at the expression, you have (1 + 3i) + (2 − 5i) and you know the rule, that first of all you have to get answers in brackets in order to so what's next? okay, all you do now is look at what stands behind each of the brackets you have (1 + 3i) while there is nothing before the symbol ( meaning that nothing in brackets changes as you "simplify" it, so you just write down : 1 + 3i now you have other part of equation to deal with witch looks like this + (2 - 5i) , okay, now let's clarify what that means when something is put in form of +(...) it means that whatever the answer may turn out to be in brackets, it will be added to sum of equation just as it is, giving you freedom to simplify it into +... since there is no multiplication or division related action going in it, giving you freedom to fullfil partof our joolly cooperation journey through equation with writting down this 1 + 3i + 2 - 5i = answer what's next, you may ask? let's define what can be simplified! take notice that digits 1 and 2 doesn't have this i thingy beside them, so you can add them one to another, getting 3 +3i - 5i = answer same goes with all numbers that have been multiplied by the same letter, so you can sum them up into -2i, therefore 3-2i =answer. hope this helps